# Context-Aware Video Instance Segmentation

### v1

@@ Abstract
In this paper, we introduce the Context-Aware Video Instance Segmentation (CAVIS), a novel framework designed to enhance instance association by integrating contextual information adjacent to each object. To efficiently extract and leverage this information, we propose the Context-Aware Instance Tracker (CAIT), which merges contextual data surrounding the instances with the core instance features to improve tracking accuracy. Additionally, we introduce the Prototypical Cross-frame Contrastive (PCC) loss, which ensures consistency in object-level features across frames, thereby significantly enhancing instance matching accuracy. CAVIS demonstrates superior performance over state-of-the-art methods on all benchmark datasets in video instance segmentation (VIS) and video panoptic segmentation (VPS). Notably, our method excels on the OVIS dataset, which is known for its particularly challenging videos. Project page: this https URL
@@ 1 Introduction
Video Instance Segmentation (VIS) is a crucial task that involves segmenting and identifying individual objects within video sequences, applicable in a variety of fields including video understanding, autonomous driving, and video editing (Yang et al., 2019). VIS has seen considerable advancements, with developments in both online methods (Yang et al., 2019; Cao et al., 2020; Yang et al., 2021b; Huang et al., 2022; Wu et al., 2022c; Ying et al., 2023; Kim et al., 2024), which process videos frame-by-frame to adapt in real-time, and offline methods (Wang et al., 2021; Hwang et al., 2021; Wu et al., 2022b; Cheng et al., 2021a; Heo et al., 2022), which analyze entire videos to understand inter-frame dependencies thoroughly.
@@
Recent advances have brought robust query-based segmentation architectures (Cheng et al., 2021b; 2022), designed to detect instance centers and cluster pixels into instance-specific groups within images. These modern VIS approaches increasingly employ instance center associations across frames to improve tracking accuracy. Techniques such as contrastive learning (Wu et al., 2022c; Ying et al., 2023) and transformer-based trackers (Heo et al., 2023; Zhang et al., 2023a) leverage the similarities between instance centers for consistent identification of objects across frames. However, challenges persist in scenarios with significant occlusions or when multiple similar objects are present, leading to potential tracking inaccuracies as shown in the top row of Fig. 1. While some strategies (Heo et al., 2022; 2023) attempt to use instance features for tracking across segments or entire videos, difficulties remain.
@@
To address this issue, we propose Context-Aware Video Instance Segmentation (CAVIS), a novel framework designed to improve object identification by incorporating contextual information surrounding each instance into the tracking process. This approach draws from insights in neuroscience and cognitive science (Bar, 2004; Oliva & Torralba, 2007), emphasizing the importance of contextual cues in human perception for deciphering complex scenes and resolving visual ambiguities. An example of the practical application of this principle is shown in Fig. 1, where recognizing a person is riding a bicycle, rather than just identifying the bicycle, greatly enhances the accuracy of object identification.
@@
To achieve this, we design the Context-Aware Instance Tracker (CAIT), featuring advanced modules for extracting and matching context-aware instance features. The context-aware instance feature extractor combines the contextual information at the object’s boundary with the core features of each instance. Then, we incorporate these context-aware instance features into a transformer-based tracking architecture (Wu et al., 2022b; Heo et al., 2023; Zhang et al., 2023a), enhanced by our novel context-aware cross-attention mechanism. This adjustment allows for the precise utilization of detailed contextual nuances within each scene.
@@
Figure 1: Importance of contextual information. Comparative results showing the state-of-the-art model (Zhang et al., 2023b) (Top) and CAVIS (Bottom). The frame on the left precedes the right by four frames, during which an occlusion takes place. The standard model, lacking contextual data, fails to consistently track the same bicycle post-occlusion, while CAVIS effectively maintains accurate instance tracking.
@@
Furthermore, we introduce the Prototypical Cross-frame Contrastive (PCC) loss to ensure feature consistency across frames. Utilizing a query-based segmenter, PCC loss correlates each pixel with an object by comparing core instance features against pixel embeddings, enhancing regional consistency within frames and establishing an intra-frame constraint reflective of the feature map’s similarity. This method crucially maps inter-frame pixel relationships, bolstering object association—a vital aspect for contrastive learning that differentiates similar from dissimilar objects across frames. However, maintaining individual pixel consistency is memory-intensive. To address this, PCC loss strategically constructs instance-wise prototypes to maintain frame-to-frame consistency, thereby boosting training efficiency and ensuring robust performance in dynamic environments.
@@
Our extensive testing shows that CAVIS significantly outperforms existing state-of-the-art methods across major video segmentation benchmarks, including YTVIS19 (Yang et al., 2019), YTVIS21 (Yang et al., 2021a), OVIS (Qi et al., 2022), and VIPSeg (Miao et al., 2021), particularly excelling on OVIS dataset that include complex video sequences. Our contributions to the field are manifold and can be summarized as follows:
  • 1.
    We present Context-Aware Instance Tracker (CAIT), a novel framework designed to extract context-aware instance features and utilize them for enhanced instance matching.
  • 2.
    We propose a Prototypical Cross-frame Contrative (PCC) loss that enhances the learning of instance matching by ensuring consistency in object-level features across frames.
  • 3.
    Our model demonstrates robustness in challenging videos environments, establishing state-of-the-art performance in Video Instance Segmentation and Video Panoptic Segmentation.
@@ 2 Related Works
Video Instance Segmentation. VIS methods learn to associate features frame-to-frame based on instance segmentation architectures. The pioneering MaskTrack R-CNN (Yang et al., 2019) integrates a tracking head into Mask R-CNN (He et al., 2017), utilizing heuristic cues for instance association Following advancements include SipMask (Cao et al., 2020) and CrossVIS (Yang et al., 2021b), which enhance temporal links through cross-frame learning. IDOL(Wu et al., 2022c) a contrastive learning approach with query-based architectures (Zhu et al., 2020), boosting online method performance. Conversely, offline approaches like VisTR (Wang et al., 2021) and Seqformer (Wu et al., 2022b) use the entire video for mask trajectory predictions, with VisTR applying DETR (Carion et al., 2020) at the clip level and Seqformer aggregating temporal information via inter-frame queries. Innovations like IFC (Hwang et al., 2021) and TeViT (Yang et al., 2022) improve efficiency by adjusting attention mechanisms within transformer architectures.
Advancements in Query-based Networks. Strong query-based segmentation networks have become prevalent in current VIS methods, with many relying on Mask2Former (Cheng et al., 2022) as their foundation. MinVIS (Huang et al., 2022) achieves tracking through simple post-processing based on cosine similarity between object features, without video learning. VITA (Heo et al., 2022) temporally associates frame-level queries to find instance prototypes within a video. GenVIS (Heo et al., 2023) adopts object association approach of VITA and designs a tracking network at the sub-clip level. Inspired by SimCLR (Chen et al., 2020), CTVIS (Ying et al., 2023) utilizes contrastive learning with a larger number of frames for comprehensive frame association. DVIS (Zhang et al., 2023a) introduces a decoupled framework for VIS, dividing it into segmentation, tracking, and refinement tasks, thereby enabling efficient and effective learning.
Object Tracking with Additional Cues. Tracking methodologies have been developed across various domains, including video object segmentation (VOS) (Xu et al., 2018; Oh et al., 2019; Cheng et al., 2021c), multiple object tracking (MOT) (Milan et al., 2016; Bergmann et al., 2019; Zhou et al., 2020; Zhang et al., 2021), and VIS. Despite advancements, many challenging cases persist, prompting research into object association with supplementary data. Early approaches leverage spatial-temporal information such as geometric relation between adjacent frames (Tang et al., 2017) and aggregated object features of previous frames (Xu et al., 2019). BeyondPixel (Sharma et al., 2018) improves inter-frame object matching by proposing a new cost that captures 3D pose and shape based on monocular geometry. BATMAN (Yu et al., 2022) combines optical flow and object query features to encode motion and appearance information into bilateral space. (Choudhuri et al., 2023) introduces a context-aware relative query which reflects global and local context by aggregating multi-level image features from consecutive frames.
@@ 3 Preliminary
This section offers a concise introduction to the fundamentals of a query-based instance segmentation pipeline and outlines a VIS approach that incorporates contrastive learning (Wu et al., 2022c; Li et al., 2023b; Ying et al., 2023; Li et al., 2023c).
@@ 3.1 Query-based instance segmentation
Modern VIS methods adopt a query-based instance segmentation pipeline (Cheng et al., 2021b; 2022), including three main components: a backbone encoder, a pixel decoder, and a transformer decoder. The backbone encoder and pixel decoder are responsible for extracting multi-scale feature maps from the input image. The transformer decoder employs object queries—sequences of latent vectors—as initial guesses for object centers and utilizes these features to generate object-level features. These queries undergo refinement through multiple transformer blocks via a cross-attention mechanism between the object queries and the feature maps. The refined instance features are then used for classification and segmentation tasks through respective prediction heads. Typically, the number of object queries, $N$, exceeds the actual number of objects, $N_{GT}$, present in the image. Traditionally, the process involves finding a permutation of $N$ elements, $\sigma\in\mathfrak{S}_{N}$, that optimally assigns the prediction set $\{\hat{y}_{i}\}_{i=1}^{N}$ to maximize total similarity to the ground truth (GT) set $\{y_{i}\}_{i=1}^{N_{GT}}$. This is achieved by minimizing a pair-wise matching cost $\mathcal{L}_{\text{Match}}$, as defined in (Cheng et al., 2021a):
@@ 3.2 Contrastive learning for VIS
In query-based architectures, the order of instance features effectively serves as the identity of each object. By aligning the sequence of instance features across frames, we can facilitate object tracking. Since instance features represent specific objects, inter-frame feature association is used for this alignment. To enhance the robustness of instance features for matching objects between frames, the following contrastive loss is integrated within the VIS framework (Wu et al., 2022c):
$\mathcal{L}_{\text{Emb}}(v_{t})$ $=-\log\frac{\sum_{k^{+}\in\mathbf{K}^{+}_{{v_{t}}}}\exp\left(v_{t}\cdot k^{+}\right)}{\sum_{k^{+}\in\mathbf{K}^{+}_{{v_{t}}}}\exp\left(v_{t}\cdot k^{+}\right)+\sum_{k^{-}\in\mathbf{K}^{-}_{{v_{t}}}}\exp\left(v_{t}\cdot k^{-}\right)},$ (4)
$=\log\left[1+\sum_{k^{+}\in\mathbf{K}^{+}_{{v_{t}}}}\sum_{k^{-}\in\mathbf{K}^{-}_{{v_{t}}}}\exp\left(v_{t}\cdot k^{-}-v_{t}\cdot k^{+}\right)\right],~{}~{}\forall t\in\{1,...,T\},$
@@ 4 Method
This section describes our Context-Aware Video Instance Segmentation (CAVIS) whose overall pipeline is illustrated in Fig. 2. Our CAVIS consists of two key components: Context-Aware Instance Tracker (CAIT) and Prototypical Cross-frame Contrastive (PCC) loss, which are detailed in Sec. 4.1 and Sec. 4.2, respectively. We describe the training losses for each network in Sec. 4.3.
@@ 4.1.1 Context-aware feature extraction
Following the method outlined in previous VIS studies (Huang et al., 2022; Heo et al., 2022; 2023; Ying et al., 2023; Zhang et al., 2023a), we employ Mask2Former (Cheng et al., 2022) as our segmentation network $\mathcal{S}$. This framework ingests a series of input frames $\{I_{t}\}_{t=1}^{T}$, with $T$ denoting the total number of frames. It extracts feature maps $F$, identifies instance features $\hat{Q}$, and computes both classification scores $O$ and generates segmentation masks $M$ as follows:
$\left\{F_{t},\hat{Q}_{t},O_{t},M_{t}\right\}_{t=1}^{T}$ $=\mathcal{S}\left(\left\{I_{t}\right\}_{t=1}^{T}\right),$ (5)
$F_{t}\in\mathbb{R}^{H\times W\times C},~{}\hat{Q}_{t}\in\mathbb{R}^{N\times C},$ $~{}O_{t}\in\mathbb{R}^{N\times K},~{}M_{t}\in\mathbb{R}^{N\times H\times W},$
@@
$\begin{split}\tilde{Q}^{n}_{t}=&\frac{\sum_{h=1}^{H}\sum_{w=1}^{W}\bar{F}^{\{h,w\}}_{t}*\mathbbm{1}\left(\acute{M}_{t}^{\{n,h,w\}}>0\right)}{\sum_{h=1}^{H}\sum_{w=1}^{W}\mathbbm{1}\left(\acute{M}_{t}^{\{n,h,w\}}>0\right)},~{}~{}\forall n=\{1,...,N\},\\ &\text{where}~{}\bar{F}=\text{Avg}(F),~{}\acute{M}=\text{Lap}(M),\end{split}$ (6)
where $\text{Avg}(\cdot)$ denotes an average filtering process, $\text{Lap}(\cdot)$ signifies the application of a Laplacian filter, and $\mathbbm{1}(\cdot)$ is the indicator function that tests for the presence of the object within the filtered mask. We employ the average filter specifically configured with a $9\times 9$ kernel size.
Finally, we combine the core and surrounding features to further enhance instance representations. The context-aware instance feature $Q^{n}_{t}\in\mathbb{R}^{C}$ is generated by concatenating the core instance feature $\hat{Q}^{n}_{t}$ and the instance surrounding feature $\tilde{Q}^{n}_{t}$, and subsequently processing this combined feature through a multi-layer perceptron (MLP) as follows:
@@ 4.1.2 Context-aware instance matching
We introduce our context-aware tracking network $\mathcal{T}$, which employs a transformer-based tracking network (Heo et al., 2023; Zhang et al., 2023a) to learn associations across adjacent frames. The network comprises six transformer blocks, each featuring cross-attention, self-attention, and feed-forward layers. The conventional cross-attention mechanism, denoted as $\text{Attn}(Q,K,V)$, traditionally aligns the current unordered instance features, $\hat{Q}_{t}$, (serving as both the key and value), with the ordered instance features from the previous frame, $\hat{Q}^{*}_{t-1}$, (used as the query). To enhance accuracy, our model employs context-aware instance features, ${Q}^{*}_{t-1}$ and ${Q}_{t}$, as the query and key, respectively while we still use the instance features $\hat{Q}_{t}$ as value. This modification leads to a context-aware cross-attention mechanism, formulated as:
@@
where $C$ is the channel dimensions of the feature maps. The aligned context-aware instance features $Q_{t}^{*}$ is built by concatenating the aligned instance features $\hat{Q}_{t}^{*}$ and the aligned instance surrounding features $\tilde{Q}_{t}^{*}$ same as in Eq. (7). We obtain the aligned instance surrounding features $\tilde{Q}_{t}^{*}$ by using Hungarian matching algorithm (Kuhn, 1955) on cosine similarity between $\hat{Q}_{t}$ and $\hat{Q}_{t}^{*}$ as follows:
$\begin{split}\tilde{Q}_{t}^{*^{\sigma_{H}(n)}}=\tilde{Q}_{t}^{n},~{}\forall n=\{1,...,N\},~{}\text{where}~{}\sigma_{H}=~{}\text{Hungarian}(\hat{Q}_{t}^{*},\hat{Q}_{t}),~{}\sigma_{H}\in\mathbb{R}^{N}.\end{split}$ (10)

### v2

# CAVIS: Context-Aware Video Instance Segmentation

@@ Abstract
In this paper, we introduce the Context-Aware Video Instance Segmentation (CAVIS), a novel framework designed to enhance instance association by integrating contextual information adjacent to each object. To efficiently extract and leverage this information, we propose the Context-Aware Instance Tracker (CAIT), which merges contextual data surrounding the instances with the core instance features to improve tracking accuracy. Additionally, we design the Prototypical Cross-frame Contrastive (PCC) loss, which ensures consistency in object-level features across frames, thereby significantly enhancing matching accuracy. CAVIS demonstrates superior performance over state-of-the-art methods on all benchmark datasets in video instance segmentation (VIS) and video panoptic segmentation (VPS). Notably, our method excels on the OVIS dataset, known for its particularly challenging videos. Project page: this https URL
@@ 1 Introduction
Video Instance Segmentation (VIS) is a crucial task that involves segmenting and identifying individual objects within video sequences, applicable in a variety of fields including video understanding, autonomous driving, and video editing [51]. VIS has seen considerable advancements, with developments in both online methods [51, 4, 53, 19, 47, 55, 22], which process videos frame-by-frame to adapt in real-time, and offline methods [44, 20, 46, 8, 17], which analyze entire videos to understand inter-frame dependencies thoroughly.
@@
Recent advances have brought robust query-based segmentation architectures [9, 10], designed to detect instance centers and cluster pixels into instance-specific groups within images. These modern VIS approaches increasingly employ instance center associations across frames to improve tracking accuracy. Techniques such as contrastive learning [47, 55] and transformer-based trackers [18, 57] leverage the similarities between instance centers for consistent identification of objects across frames. However, challenges persist in scenarios with significant occlusions or when multiple similar objects are present, leading to potential tracking inaccuracies as shown in the top row of Fig. 1. While some strategies [17, 18] attempt to use instance features for tracking across segments or entire videos, difficulties remain.
@@
To address this issue, we propose Context-Aware Video Instance Segmentation (CAVIS), a novel framework designed to improve object identification by incorporating contextual information surrounding each instance into the tracking process. This approach draws from insights in neuroscience and cognitive science [2, 36], emphasizing the importance of contextual cues in human perception for deciphering complex scenes and resolving visual ambiguities. An example of the practical application of this principle is shown in Fig. 1, where recognizing a person is riding a bicycle, rather than just identifying the bicycle, greatly enhances the accuracy of object identification.
@@
To achieve this, we design the Context-Aware Instance Tracker (CAIT), featuring advanced modules for extracting and matching context-aware instance features. The context-aware instance feature extractor combines the contextual information at the object’s boundary with the core features of each instance. Then, we incorporate these context-aware instance features into a transformer-based tracking architecture [46, 18, 57], enhanced by our context-aware cross-attention mechanism. This adjustment allows for the precise utilization of detailed contextual nuances within each scene.
@@
Figure 1: Importance of contextual information. Comparative results showing the state-of-the-art model [58] (Top) and CAVIS (Bottom). The frame on the left precedes the right by four frames, during which an occlusion takes place. The standard model, lacking contextual data, fails to consistently track the same bicycle post-occlusion, while CAVIS effectively maintains accurate instance tracking.
@@
Furthermore, inspired by a recent study [23], we design the Prototypical Cross-frame Contrastive (PCC) loss, specifically designed to enhance feature consistency across frames. PCC loss correlates each pixel with an object by comparing core instance features against pixel embeddings, enhancing regional consistency within frames and establishing an intra-frame constraint reflective of the feature map’s similarity. This method leverages the close relationship between object features and high-level feature maps, where their similarity guides mask predictions. By constructing instance-wise prototypes, PCC loss maintains frame-to-frame consistency, improving training efficiency and ensuring robust performance in dynamic environments.
@@
Our extensive testing shows that CAVIS significantly outperforms existing state-of-the-art methods across major video segmentation benchmarks, including YTVIS19 [51], YTVIS21 [52], OVIS [38], and VIPSeg [32], particularly excelling on OVIS dataset that includes complex videos.
@@ 2 Related Works
Video Instance Segmentation. VIS methods learn to associate features frame-to-frame based on instance segmentation architectures. The pioneering MaskTrack R-CNN [51] integrates a tracking head into Mask R-CNN [16], utilizing heuristic cues for instance association Following advancements include SipMask [4] and CrossVIS [53], which enhance temporal links through cross-frame learning. IDOL[47] a contrastive learning approach with query-based architectures [61], boosting online method performance. Conversely, offline approaches like VisTR [44] and Seqformer [46] use the entire video for mask trajectory predictions, with VisTR applying DETR [5] at the clip level and Seqformer aggregating temporal information via inter-frame queries. Innovations like IFC [20] and TeViT [54] improve efficiency by adjusting attention mechanisms within transformer architectures.
Advancements in Query-based Networks. Strong query-based segmentation networks have become prevalent in current VIS methods, with many relying on Mask2Former [10] as their foundation. MinVIS [19] achieves tracking through simple post-processing based on cosine similarity between object features, without video learning. VITA [17] temporally associates frame-level queries to find instance prototypes within a video. GenVIS [18] adopts object association approach of VITA and designs a tracking network at the sub-clip level. Inspired by SimCLR [6], CTVIS [55] utilizes contrastive learning with a larger number of frames for comprehensive frame association. DVIS [57] introduces a decoupled framework for VIS, dividing it into segmentation, tracking, and refinement tasks, thereby enabling efficient and effective learning.
Object Tracking with Additional Cues. Tracking methodologies have been developed across various domains, including video object segmentation (VOS) [50, 35, 11], multiple object tracking (MOT) [34, 3, 60, 59], and VIS. Despite advancements, many challenging cases persist, prompting research into object association with supplementary data. Early approaches leverage spatial-temporal information such as geometric relation between adjacent frames [42] and aggregated object features of previous frames [49]. BeyondPixel [40] improves inter-frame object matching by proposing a new cost that captures 3D pose and shape based on monocular geometry. BATMAN [56] combines optical flow and object query features to encode motion and appearance information into bilateral space. CAROQ [12] employs a context feature defined as a memory bank of multi-level image features extracted by a pixel decoder. However, such a full-context-based approach can lead to increased complexity and memory limitations. In contrast, our method takes a memory-efficient approach by focusing on the surrounding features of each object during tracking, enabling effective object matching.
@@ 3 Preliminary
This section offers a concise introduction to the fundamentals of a query-based segmentation pipeline and outlines a VIS approach that incorporates contrastive loss [47, 26, 55, 28].
@@ 3.1 Query-based instance segmentation
Modern VIS methods adopt a query-based instance segmentation pipeline [9, 10], including three main components: a backbone encoder, a pixel decoder, and a transformer decoder. The backbone encoder and pixel decoder are responsible for extracting multi-scale feature maps from the input image. The transformer decoder employs object queries—sequences of latent vectors—as initial guesses for object centers and utilizes these features to generate object-level features. These queries undergo refinement through multiple transformer blocks via a cross-attention mechanism between the object queries and the feature maps. The refined instance features are then used for classification and segmentation tasks through respective prediction heads. Typically, the number of object queries, $N$, exceeds the actual number of objects, $N_{GT}$, present in the image. Traditionally, the process involves finding a permutation of $N$ elements, $\sigma\in\mathfrak{S}_{N}$, that optimally assigns the prediction set $\{\hat{y}_{i}\}_{i=1}^{N}$ to maximize total similarity to the ground truth (GT) set $\{y_{i}\}_{i=1}^{N_{GT}}$. This is achieved by minimizing a pair-wise matching cost $\mathcal{L}_{\text{Match}}$, as defined in [8]:
@@
Table 1: Notations used in our method.
Symbol Description Symbol Description $\hat{Q}$ : instance features $M$ : mask predictions $\tilde{Q}$ : instance surrounding features $\acute{M}$ : boundary scores processed from $M$ ${Q}$ : context-aware instance features $F$ : the last feature maps from pixel decoder ${Q}^{*}$ : aligned context-aware instance features $\bar{F}$ : feature maps processed by average filter
@@ 3.2 Contrastive learning for VIS
In query-based architectures, the order of instance features effectively serves as the identity of each object. By aligning the sequence of instance features across frames, we can facilitate object tracking. Since instance features represent specific objects, inter-frame feature association is used for this alignment. To enhance the robustness of instance features for matching objects between frames, the following contrastive loss is integrated within the VIS framework [47]:
$\mathcal{L}_{\text{Emb}}(v_{t})&=\log\left[1+\sum_{k^{+}\in\mathbf{K}^{+}_{{v_{t}}}}\sum_{k^{-}\in\mathbf{K}^{-}_{{v_{t}}}}\exp\left(v_{t}\cdot k^{-}-v_{t}\cdot k^{+}\right)\right],\end{aligned}$
@@ 4 Method
This section describes our Context-Aware Video Instance Segmentation (CAVIS) whose overall pipeline is illustrated in Fig. 2. Our CAVIS consists of two key components: Context-Aware Instance Tracker (CAIT) and Prototypical Cross-frame Contrastive (PCC) loss, which are detailed in Sec. 4.1 and Sec. 4.2, respectively. We describe the training losses for each network in Sec. 4.3. We provide a notation table in Tab. 1 for better readability.
@@ 4.1.1 Context-aware feature extraction
Following the method outlined in previous VIS studies [19, 17, 18, 55, 57], we employ Mask2Former [10] as our segmentation network $\mathcal{S}$. This framework ingests a series of input frames $\{I_{t}\}_{t=1}^{T}$, with $T$ denoting the total number of frames. It extracts feature maps $F$, identifies instance features $\hat{Q}$, and computes both classification scores $O$ and generates segmentation masks $M$ as follows:
[…]
@@
$\begin{split}\tilde{Q}^{n}_{t}=\frac{\sum_{h=1}^{H}\sum_{w=1}^{W}\bar{F}^{\{h,w\}}_{t}*\mathbbm{1}\left(\acute{M}_{t}^{\{n,h,w\}}>0\right)}{\sum_{h=1}^{H}\sum_{w=1}^{W}\mathbbm{1}\left(\acute{M}_{t}^{\{n,h,w\}}>0\right)},\\ \bar{F}=\text{Avg}(F),~{}\acute{M}=\text{Lap}(M),~{}\forall n=\{1,...,N\},\end{split}$ (6)
where $\text{Avg}(\cdot)$ denotes an average filtering process over spatial dimensions, $\text{Lap}(\cdot)$ signifies the application of a Laplacian filter, and $\mathbbm{1}(\cdot)$ is the indicator function that tests for the presence of the object within the filtered mask. We employ the average filter specifically configured with a $9\times 9$ kernel.
Finally, we combine the core and surrounding features to further enhance instance representations. The context-aware instance feature $Q^{n}_{t}\in\mathbb{R}^{C}$ is generated by concatenating the core instance feature $\hat{Q}^{n}_{t}$ and the instance surrounding feature $\tilde{Q}^{n}_{t}$ along the channel dimension, and subsequently processing this combined feature through a multi-layer perceptron (MLP) as follows:
@@
Figure 3: Qualitative comparisons of CAVIS (ours) against state-of-the-art methods: CTVIS [55] and DVIS++ [58] on the OVIS dataset.
@@ 4.1.2 Context-aware instance matching
We introduce our context-aware tracking network $\mathcal{T}$, which employs a transformer-based tracking network [18, 57] to learn associations across adjacent frames. The network comprises six transformer blocks, each featuring cross-attention, self-attention, and feed-forward layers. The conventional cross-attention mechanism, denoted as $\text{Attn}(Q,K,V)$, traditionally aligns the current unordered instance features, $\hat{Q}_{t}$, (serving as both the key and value), with the ordered instance features from the previous frame, $\hat{Q}^{*}_{t-1}$, (used as the query). To enhance accuracy, our model employs context-aware instance features, ${Q}^{*}_{t-1}$ and ${Q}_{t}$, as the query and key, respectively while we still use the instance features $\hat{Q}_{t}$ as value. This modification leads to a context-aware cross-attention mechanism, formulated as:
@@
where $C$ is the channel dimensions of the feature maps. The aligned context-aware instance features $Q_{t}^{*}$ is built by concatenating the aligned instance features $\hat{Q}_{t}^{*}$ and the aligned instance surrounding features $\tilde{Q}_{t}^{*}$ same as in Eq. (7). We obtain the aligned instance surrounding features $\tilde{Q}_{t}^{*}$ by using Hungarian matching algorithm [24] on cosine similarity between $\hat{Q}_{t}$ and $\hat{Q}_{t}^{*}$ as follows:
$\begin{gathered}\tilde{Q}_{t}^{*^{\sigma_{H}(n)}}=\tilde{Q}_{t}^{n},~{}\forall n=\{1,...,N\},\\ ~{}\text{where}~{}\sigma_{H}=~{}\text{Hungarian}(\hat{Q}_{t}^{*},\hat{Q}_{t}),~{}\sigma_{H}\in\mathbb{R}^{N}.\end{gathered}$ (12)
@@
Table 2: Comparisons on the validation sets of YouTube-VIS 2019, 2021, and OVIS datasets. The best and second-best scores are highlighted in red and blue, respectively. denotes the model trained with the temporal refiner [57]. Rows in cyan indicate comparisons with a top-performing model.
Methods OVIS YTVIS19 YTVIS21 AP AP50 AP75 AR1 AR10 AP AP50 AP75 AR1 AR10 AP AP50 AP75 AR1 AR10 ResNet-50 MaskTrack R-CNN [51] 10.8 25.3 8.5 7.9 14.9 30.3 51.1 32.6 31.0 35.5 28.6 48.9 29.6 26.5 33.8 SipMask [4] 10.2 24.7 7.8 7.9 15.8 33.7 54.1 35.8 35.4 40.1 31.7 52.5 34.0 30.8 37.8 IFC [20] 13.1 27.8 11.6 9.4 23.9 41.2 65.1 44.6 42.3 49.6 35.2 55.9 37.7 32.6 42.9 CrossVIS [53] 14.9 32.7 12.1 10.3 19.8 36.3 56.8 38.9 35.6 40.7 34.2 54.4 37.9 30.4 38.2 EfficientVIS [48] - - - - - 37.9 59.7 43.0 40.3 46.6 34.0 57.5 37.3 33.8 42.5 SeqFormer [46] 15.1 31.9 13.8 10.4 27.1 47.4 69.8 51.8 45.5 54.8 40.5 62.4 43.7 36.1 48.1 VISOLO [14] 15.3 31.0 13.8 11.1 21.7 38.6 56.3 43.7 35.7 42.5 36.9 54.7 40.2 30.6 40.9 Mask2Former-VIS [8] 17.3 37.3 15.1 10.5 23.5 46.4 68.0 50.0 - - 40.6 60.9 41.8 - - VITA [17] 19.6 41.2 17.4 11.7 26.0 49.8 72.6 54.5 49.4 61.0 45.7 67.4 49.5 40.9 53.6 MinVIS [19] 25.0 45.5 24.0 13.9 29.7 47.4 69.0 52.1 45.7 55.7 44.2 66.0 48.1 39.2 51.7 CAROQ [12] 25.8 47.9 25.4 14.2 33.9 46.7 70.4 50.9 45.7 55.9 43.3 64.9 47.1 39.3 52.7 IDOL [47] 28.2 51.0 28.0 14.5 38.6 49.5 74.0 52.9 47.7 58.7 43.9 68.0 49.6 38.0 50.9 DVIS [57] 30.2 55.0 30.5 14.5 37.3 51.2 73.8 57.1 47.2 59.3 46.4 68.4 49.6 39.7 53.5 TCOVIS [25] 35.3 60.7 36.6 15.7 39.5 52.3 73.5 57.6 49.8 60.2 49.5 71.2 53.8 41.3 55.9 CTVIS [55] 35.5 60.8 34.9 16.1 41.9 55.1 78.2 59.1 51.9 63.2 50.1 73.7 54.7 41.8 59.5 GenVIS [18] 35.8 60.8 36.2 16.3 39.6 50.0 71.5 54.6 49.5 59.7 47.1 67.5 51.5 41.6 54.7 VISAGE [23] 36.2 60.3 35.3 16.1 40.3 55.1 78.1 60.6 51.0 62.3 51.6 73.8 56.1 43.6 59.3 Ours 37.6 63.4 38.2 16.5 43.5 55.7 78.3 61.7 51.5 63.3 50.5 74.1 54.9 42.6 58.5 Swin-L SeqFormer [46] - - - - - 59.3 82.1 66.4 51.7 64.4 51.8 74.6 58.2 42.8 58.1 Mask2Former-VIS [8] 25.8 46.5 24.4 13.7 32.2 60.4 84.4 67.0 - - 52.6 76.4 57.2 - - VITA [17] 27.7 51.9 24.9 14.9 33.0 63.0 86.9 67.9 56.3 68.1 57.5 80.6 61.0 47.7 62.6 CAROQ [12] 38.2 60.7 39.5 17.7 44.1 61.4 82.8 68.6 55.2 68.1 54.5 75.4 60.5 45.5 61.4 MinVIS [19] 39.4 61.5 41.3 18.1 43.3 61.6 83.3 68.6 54.8 66.6 55.3 76.6 62.0 45.9 60.8 IDOL [47] 40.0 63.1 40.5 17.6 46.4 64.3 87.5 71.0 55.6 69.1 56.1 80.8 63.5 45.0 60.1 GenVIS [18] 45.2 69.1 48.4 19.1 48.6 64.0 84.9 68.3 56.1 69.4 59.6 80.9 65.8 48.7 65.0 DVIS [57] 45.9 71.1 48.3 18.5 51.5 63.9 87.2 70.4 56.2 69.0 58.7 80.4 66.6 47.5 64.6 TCOVIS [25] 46.7 70.9 49.5 19.1 50.8 64.1 86.6 69.5 55.8 69.0 61.3 82.9 68.0 48.6 65.1 CTVIS [55] 46.9 71.5 47.5 19.1 52.1 65.6 87.7 72.2 56.5 70.4 61.2 84.0 68.8 48.0 65.8 Ours 48.6 74.0 52.5 19.5 53.3 66.0 89.5 73.3 56.8 71.4 61.1 84.1 69.2 48.2 66.3 ViT-L MinVIS [19] 42.9 65.7 45.4 19.8 46.5 65.6 85.4 72.7 57.5 70.6 59.2 79.9 66.7 47.8 64.1 DVIS++ [58] 49.6 72.5 55.0 20.8 54.6 67.7 88.8 75.3 57.9 73.7 62.3 82.7 70.2 49.5 68.0 Ours 53.2 75.9 59.1 20.9 58.2 68.9 89.3 76.2 58.3 73.6 64.6 85.6 72.5 49.5 69.3 DVIS++ [58] 53.4 78.9 58.5 21.1 58.7 68.3 90.3 76.1 57.8 73.4 63.9 86.7 71.5 48.8 69.5 Ours 57.1 82.6 63.5 21.2 61.8 69.4 90.9 77.2 58.3 74.7 65.3 87.3 73.2 49.7 70.3